OpenStudy (anonymous):

Find S15 for 1 + 1.5 + 2.25 + 3.375 + … i know the answer is 873.8, but can someone please give me an easier formula to get this answer?

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

there's just the standard formula we used in the previous question r= 1.5 \(\huge a_n=a_1*r^{n-1}\) \(\huge a_n=1*1.5^{n-1}\)

OpenStudy (bibby):

oops

OpenStudy (bibby):

are they asking for the 15th term or what

OpenStudy (anonymous):

yes

OpenStudy (bibby):

yeah, same formula as last time. how'd you get 873.8

OpenStudy (anonymous):

by dividing the number by 2 then adding the same number to the answer we got after dividing... example 2.25/2=1.125 then 1.125+2.25=3.375

OpenStudy (bibby):

ok, first of all to find the common ratio we divide a term by the term before it \(\huge \frac{1.5}{1}=\frac{2.25}{1.5}=\frac{3.375}{2.25}=1.5\)

OpenStudy (anonymous):

ok that makes it easier! thank you!

OpenStudy (bibby):

so what's our next step?

OpenStudy (anonymous):

multiplying each number by 1.5?

OpenStudy (bibby):

assuming they want the 15th term, no. we just plug into the formula

OpenStudy (anonymous):

whats n?

OpenStudy (bibby):

nth term means the term we want to find. if we want \(s_{15}\) n=15

OpenStudy (anonymous):

but the answer that gives me is 291.9

OpenStudy (bibby):

how do you know what the answer should be

OpenStudy (bibby):

that would mean we're finding the partial sum, not the n-th term

OpenStudy (bibby):

\(\huge S_n=\frac{a_1(1-r^n}{1-r}\)

OpenStudy (anonymous):

ohhh ok! that gave me the answer i needed! thank you(:

OpenStudy (bibby):

npnp p.s. your smiley face is backwards :)*